The volume of a box must be 315 cubic meter. The length of the box is to be 7 meters and the height is to be 4 meters more than the width. Find the width of the box

Answer by checkley77(12569)

volume=l*w*h
315=7*w*(w+4)
315=7w(w+4)
315=7w^2+28w
7w^2+28w-315=0
7(w^2+4w-45)=0
7(w+9)(w-5)=0
w-5=0
w=5 ans. for the width.
Proof:
315=7*5(5+4)
315=35*9
315=315